OpenStudy (anonymous):

subtract x/x-2 x+1 - 2/ x-1 x+1

OpenStudy (anonymous):

You should use paretheses becuase i can tell which parts are fractions

OpenStudy (anonymous):

(x/(x-2)(x+1)- 2/ (x-1)(x+1)

OpenStudy (lalaly):

\[=(\frac{x}{(x-2)(x+1)} \times \frac{x-1}{x-1})-( \frac{2}{(x-1)(x+1)} \times \frac{x-2}{x-2})\]

OpenStudy (lalaly):

\[=\frac{x(x-1)-2(x-2)}{(x-1)(x-2)(x+1)}\]

OpenStudy (lalaly):

\[=\frac{x^2-3x+4}{(x-1)(x-2)(x+1)}\]

OpenStudy (anonymous):

thanks sweetheart

OpenStudy (lalaly):

:D:D